# Deep Positive-Unlabeled Anomaly Detection for Contaminated Unlabeled Data

### v1

@@ 5.5 Hyperparameter Sensitivity
Figure 2: Relationship between the anomaly detection performance and the hyperparameter $\alpha$ of the PUAE on each dataset. We used class 0 as unseen anomaly, class 1 as normal, and the remaining classes as seen anomalies. We plotted the AUROC of the ABC by a dashed line for comparison. The semi-transparent area represents standard deviations. The anomaly detection performance is improved by tuning $\alpha$.
Our approach has the hyperparameter $\alpha$, which represents the probability of anomaly occurrence. In the above experiments, we set it to $\alpha=0.1$ since the 10% of the training dataset is seen anomalies. Although $\alpha$ can be estimated from the unlabeled and anomaly training data in conventional PU learning approaches [29, 44, 17, 6], its nature in our approach has not yet been clarified.
Finally, we evaluate the sensitivity of $\alpha$ in our approach. Figure 2 shows the relationship between the anomaly detection performance and the hyperparameter $\alpha$ of the PUAE on each dataset. We used class 0 as unseen anomaly, class 1 as normal, and the remaining classes as seen anomalies on each dataset.
It can be seen that the performance is improved by tuning $\alpha$. Note that the performance is better when $\alpha$ is set to a larger value than $0.1$, which is the true probability of anomaly occurrence. The possible reasons for this are as follows. In the objective function of our approach (Eq. (13)), the larger $\alpha$, the greater the effect of maximizing the anomaly scores for the anomaly data. In the case of conventional PU learning for binary classification, if $\alpha$ is set unnecessarily large, the decision boundary of the classifier will be biased. On the other hand, since the PUAE is based on the autoencoder, it is possible to maintain the ability to reconstruct normal data points even if $\alpha$ becomes large to some extent. Therefore, for improving the anomaly detection performance, it is better to increase $\alpha$ so that the PUAE fails to reconstruct anomalies. At present, we should determine $\alpha$ based on the validation dataset. The analysis regarding the optimal $\alpha$ is part of our future work.
The above experiments and discussion show that our approach is effective on various datasets, and that we can further improve its performance by selecting the appropriate base detector and tuning the hyperparameter.
6 Conclusion and Limitations
Although unlabeled data are assumed to be mostly normal in anomaly detection, they are often contaminated with anomalies in practice, which prevents us from improving the detection performance. To solve this, we propose the positive-unlabeled autoencoder, the semi-supervised anomaly detection approach based on PU learning and the anomaly detector such as the AE and the DeepSVDD. Our approach enables us to approximate the anomaly scores for normal data with the unlabeled and anomaly data. Therefore, without the labeled normal data, we can train the anomaly detector to minimize the anomaly scores for normal data, and to maximize those for the anomaly data. Our approach achieves better detection performance than existing approaches on various datasets.
Our approach has the following two limitations. First, our approach assume that the unlabeled data do not contain unseen anomalies. We can handle such unseen anomalies by using the robust anomaly detector for our approach, such as the LOE. Second, our approach has a hyperparameter that represents the probability of anomaly occurrence. Although this hyperparameter can be estimated by conventional PU learning approaches, we can further improve the detection performance by tuning it, as discussed in Section 5.5. Analyzing this hyperparameter is our important future work.
@@ Appendix A Dataset Example
Appendix B Additional Results
Table 2: Comparison of anomaly detection performance for seen anomalies.
@@ Appendix B Additional Results
IF [27] 0.814 $\pm$ 0.095 0.913 $\pm$ 0.056 0.514 $\pm$ 0.015 0.566 $\pm$ 0.098
AE [16] 0.882 $\pm$ 0.062 0.822 $\pm$ 0.097 0.597 $\pm$ 0.024 0.545 $\pm$ 0.125
DeepSVDD [46] 0.758 $\pm$ 0.061 0.661 $\pm$ 0.071 0.514 $\pm$ 0.016 0.556 $\pm$ 0.033
LOE [43] 0.893 $\pm$ 0.056 0.861 $\pm$ 0.105 0.593 $\pm$ 0.026 0.547 $\pm$ 0.125
ABC [59] 0.885 $\pm$ 0.061 0.825 $\pm$ 0.097 0.598 $\pm$ 0.024 0.546 $\pm$ 0.125
DeepSAD [47] 0.793 $\pm$ 0.055 0.690 $\pm$ 0.072 0.516 $\pm$ 0.016 0.559 $\pm$ 0.036
SOEL [25] 0.937 $\pm$ 0.039 0.899 $\pm$ 0.077 0.600 $\pm$ 0.029 0.549 $\pm$ 0.123
PU [22] 0.951 $\pm$ 0.063 0.817 $\pm$ 0.182 0.503 $\pm$ 0.013 0.509 $\pm$ 0.069
PUAE 0.970 $\pm$ 0.021 0.943 $\pm$ 0.041 0.608 $\pm$ 0.027 0.568 $\pm$ 0.115
PUSVDD 0.929 $\pm$ 0.028 0.891 $\pm$ 0.054 0.540 $\pm$ 0.032 0.576 $\pm$ 0.039
Table 3: Comparison of anomaly detection performance for unseen anomalies.
@@
IF [27] 0.954 $\pm$ 0.033 0.917 $\pm$ 0.112 0.495 $\pm$ 0.012 0.533 $\pm$ 0.077
AE [16] 0.996 $\pm$ 0.003 0.847 $\pm$ 0.113 0.607 $\pm$ 0.037 0.431 $\pm$ 0.104
DeepSVDD [46] 0.806 $\pm$ 0.065 0.712 $\pm$ 0.081 0.515 $\pm$ 0.015 0.558 $\pm$ 0.031
LOE [43] 0.996 $\pm$ 0.003 0.873 $\pm$ 0.119 0.606 $\pm$ 0.036 0.429 $\pm$ 0.104
ABC [59] 0.996 $\pm$ 0.004 0.848 $\pm$ 0.113 0.607 $\pm$ 0.038 0.432 $\pm$ 0.104
DeepSAD [47] 0.815 $\pm$ 0.065 0.723 $\pm$ 0.084 0.515 $\pm$ 0.014 0.560 $\pm$ 0.030
SOEL [25] 0.996 $\pm$ 0.003 0.887 $\pm$ 0.119 0.608 $\pm$ 0.041 0.432 $\pm$ 0.102
PU [22] 0.877 $\pm$ 0.181 0.774 $\pm$ 0.223 0.491 $\pm$ 0.010 0.316 $\pm$ 0.047
PUAE 0.978 $\pm$ 0.015 0.915 $\pm$ 0.118 0.612 $\pm$ 0.041 0.452 $\pm$ 0.097
PUSVDD 0.868 $\pm$ 0.074 0.846 $\pm$ 0.125 0.531 $\pm$ 0.027 0.565 $\pm$ 0.036
As the additional experiments, we evaluate the anomaly detection performance for seen and unseen anomalies on each dataset. The setups are the same as those described in Section 5. For the test data, we sampled normal data points, seen anomalies, and unseen anomalies from the test dataset up to 1,000 samples each.
Tables 2 and 3 show the anomaly detection performance for seen and unseen anomalies, respectively. We showed the average of the AUROC scores for all normal classes. We used bold to highlight the best results and statistically non-different results according to a pair-wise $t$-test. We used 5% as the p-value.
For seen anomalies, the PUAE achieved the best performance among all detectors. This shows the effectiveness of our approach, which is robust to the contaminated unlabeled data according to PU learning. For unseen anomalies, the PUAE achieved the best performance on FashionMNIST and SVHN, and the PUSVDD achieved the best performance on CIFAR10. This indicates that we may be able to improve the detection performance for unseen anomalies by using seen anomalies.
These results also show the difference between the conventional PU learning and our approach. The PU achieved the poor detection performance for unseen anomalies. This is because that it sets the decision boundary between normal data points and seen anomalies. On the other hand, our approach can detect unseen anomalies since it is based on the anomaly detector.

### v2

@@ 5.5 Hyperparameter Sensitivity
Our approach and the SOEL have the hyperparameter $\alpha$, which represents the probability of anomaly occurrence. In the above experiments, we set it to $\alpha=0.1$ since the 10% of the training dataset is anomalies. Finally, we evaluate the sensitivity of $\alpha$. Figure 2 shows the relationship between the anomaly detection performance and the hyperparameter $\alpha$ of the PUSVDD and the SOEL on each dataset.
@@
Figure 2: Relationship between the anomaly detection performance and the hyperparameter $\alpha$ of the PUSVDD and the SOEL on each dataset. We used class 0 as unseen anomaly, class 1 as normal, and the remaining classes as seen anomalies. The semi-transparent area represents standard deviations.
In most datasets, the PUSVDD and the SOEL achieve the best performance around $\alpha=0.1$. This indicates that, as with conventional PU learning, it is important to set $\alpha$ accurately. Note that $\alpha$ can be estimated from the unlabeled and anomaly training data using conventional PU learning approaches (Menon et al., 2015; Ramaswamy et al., 2016; Jain et al., 2016; Christoffel et al., 2016).
Compared to the SOEL, the PUSVDD is more robust to variations in $\alpha$. In other words, even if $\alpha$ deviates from the true value, the PUSVDD maintains relatively stable performance. This is because $\alpha$ in the SOEL is closely related to the number of anomalies within the unlabeled data. If $\alpha$ deviates from the true value, normal data may be incorrectly treated as anomalies, or vice versa.
On the other hand, since $\alpha$ in the PUSVDD only adjusts the weight of the loss function, it is expected to be relatively robust to deviations in $\alpha$.
6 Conclusion
Although most unlabeled data are assumed to be normal in semi-supervised anomaly detection, they are often contaminated with anomalies in practice, which prevents us from improving the detection performance. To solve this, we propose the deep positive-unlabeled anomaly detection framework, which integrates PU learning with deep anomaly detection models such as the AE and the DeepSVDD. Our approach enables us to approximate the anomaly scores for normal data with the unlabeled data and labeled anomaly data. Therefore, without the labeled normal data, we can train the anomaly detector to minimize the anomaly scores for normal data, and to maximize those for the anomaly data. Our approach achieves better detection performance than existing approaches on various image datasets. In the future, we will extend our approach to time-series data.
@@ Appendix A Dataset Example
Appendix B Anomaly Detection Performance with Various Numbers of Unlabeled Anomalies
(a) MNIST
(b) FashionMNIST
(c) SVHN
(d) CIFAR10
Figure 4: Comparison of anomaly detection performance between the PUSVDD and the SOEL with various numbers of unlabeled anomalies on each dataset. We used class 0 as unseen anomaly, class 1 as normal, and the remaining classes as seen anomalies. The semi-transparent area represents standard deviations.
Figure 4 shows the anomaly detection performance with various numbers of unlabeled anomalies. The hyperparameter $\alpha$ was set to its true value for each case. Our PUSVDD achieved equal to or better performance than the SOEL.
Appendix C Anomaly Detection Performance for Seen and Unseen Anomalies
Tables 3, 4, 5 and 6 show the anomaly detection performance for seen and unseen anomalies, respectively. We showed the average of the AUROC scores for all normal classes. We used bold to highlight the best results and statistically non-different results according to a pair-wise $t$-test. We used 5% as the p-value. For seen anomalies, the PUSVDD achieved the best performance among all approaches. This shows the effectiveness of our approach, which is robust to the contaminated unlabeled data according to PU learning. For unseen anomalies, although the performance is highly dataset-dependent, the PUSVDD generally performs well. This indicates that we may be able to improve the detection performance for unseen anomalies by using seen anomalies. These results also show the difference between the conventional PU learning and our approach. The PU achieved the poor detection performance for unseen anomalies. This is because that it sets the decision boundary between normal data points and seen anomalies. On the other hand, our approach can detect unseen anomalies since it is based on the anomaly detector.
Table 3: Comparison of anomaly detection performance for seen anomalies on image datasets.
@@ Appendix C Anomaly Detection Performance for Seen and Unseen Anomalies
IF [Liu et al., 2008] 0.814 $\pm$ 0.096 0.915 $\pm$ 0.053 0.510 $\pm$ 0.016 0.585 $\pm$ 0.106
AE [Hinton and Salakhutdinov, 2006] 0.843 $\pm$ 0.073 0.840 $\pm$ 0.092 0.563 $\pm$ 0.039 0.573 $\pm$ 0.131
DeepSVDD [Ruff et al., 2018] 0.925 $\pm$ 0.056 0.942 $\pm$ 0.040 0.600 $\pm$ 0.033 0.694 $\pm$ 0.069
LOE [Qiu et al., 2022] 0.942 $\pm$ 0.038 0.940 $\pm$ 0.042 0.645 $\pm$ 0.055 0.713 $\pm$ 0.077
ABC [Yamanaka et al., 2019] 0.852 $\pm$ 0.072 0.841 $\pm$ 0.092 0.564 $\pm$ 0.039 0.572 $\pm$ 0.131
DeepSAD [Ruff et al., 2019] 0.930 $\pm$ 0.052 0.956 $\pm$ 0.032 0.674 $\pm$ 0.031 0.716 $\pm$ 0.074
SOEL [Li et al., 2023] 0.967 $\pm$ 0.024 0.963 $\pm$ 0.028 0.751 $\pm$ 0.035 0.773 $\pm$ 0.073
PU [Kiryo et al., 2017] 0.959 $\pm$ 0.036 0.943 $\pm$ 0.057 0.678 $\pm$ 0.099 0.703 $\pm$ 0.154
PUAE 0.980 $\pm$ 0.017 0.942 $\pm$ 0.041 0.716 $\pm$ 0.053 0.695 $\pm$ 0.083
PUSVDD 0.994 $\pm$ 0.004 0.972 $\pm$ 0.029 0.787 $\pm$ 0.069 0.796 $\pm$ 0.059
Table 4: Comparison of anomaly detection performance for unseen anomalies on image datasets.
@@
IF [Liu et al., 2008] 0.955 $\pm$ 0.031 0.917 $\pm$ 0.111 0.492 $\pm$ 0.018 0.635 $\pm$ 0.095
AE [Hinton and Salakhutdinov, 2006] 0.981 $\pm$ 0.013 0.842 $\pm$ 0.130 0.560 $\pm$ 0.045 0.497 $\pm$ 0.111
DeepSVDD [Ruff et al., 2018] 0.950 $\pm$ 0.044 0.900 $\pm$ 0.143 0.564 $\pm$ 0.043 0.724 $\pm$ 0.087
LOE [Qiu et al., 2022] 0.949 $\pm$ 0.034 0.892 $\pm$ 0.152 0.602 $\pm$ 0.058 0.724 $\pm$ 0.099
ABC [Yamanaka et al., 2019] 0.980 $\pm$ 0.014 0.840 $\pm$ 0.133 0.559 $\pm$ 0.046 0.497 $\pm$ 0.109
DeepSAD [Ruff et al., 2019] 0.954 $\pm$ 0.040 0.900 $\pm$ 0.153 0.631 $\pm$ 0.044 0.735 $\pm$ 0.078
SOEL [Li et al., 2023] 0.963 $\pm$ 0.033 0.908 $\pm$ 0.135 0.702 $\pm$ 0.061 0.778 $\pm$ 0.087
PU [Kiryo et al., 2017] 0.965 $\pm$ 0.038 0.901 $\pm$ 0.134 0.684 $\pm$ 0.102 0.683 $\pm$ 0.133
PUAE 0.985 $\pm$ 0.017 0.894 $\pm$ 0.140 0.662 $\pm$ 0.078 0.639 $\pm$ 0.090
PUSVDD 0.983 $\pm$ 0.024 0.924 $\pm$ 0.134 0.708 $\pm$ 0.103 0.811 $\pm$ 0.101
Table 5: Comparison of anomaly detection performance for seen anomalies on real datasets.
CIFAR100 PathMNIST OCTMNIST TissueMNIST
IF [Liu et al., 2008] 0.577 $\pm$ 0.005 0.657 $\pm$ 0.211 0.679 $\pm$ 0.005 0.443 $\pm$ 0.189
AE [Hinton and Salakhutdinov, 2006] 0.514 $\pm$ 0.013 0.562 $\pm$ 0.253 0.808 $\pm$ 0.005 0.443 $\pm$ 0.188
DeepSVDD [Ruff et al., 2018] 0.623 $\pm$ 0.036 0.769 $\pm$ 0.139 0.702 $\pm$ 0.043 0.692 $\pm$ 0.047
LOE [Qiu et al., 2022] 0.624 $\pm$ 0.035 0.746 $\pm$ 0.167 0.750 $\pm$ 0.031 0.657 $\pm$ 0.084
ABC [Yamanaka et al., 2019] 0.516 $\pm$ 0.014 0.564 $\pm$ 0.252 0.805 $\pm$ 0.002 0.448 $\pm$ 0.187
DeepSAD [Ruff et al., 2019] 0.628 $\pm$ 0.042 0.772 $\pm$ 0.165 0.798 $\pm$ 0.032 0.715 $\pm$ 0.040
SOEL [Li et al., 2023] 0.696 $\pm$ 0.020 0.806 $\pm$ 0.142 0.825 $\pm$ 0.017 0.739 $\pm$ 0.044
PU [Kiryo et al., 2017] 0.630 $\pm$ 0.023 0.847 $\pm$ 0.097 0.565 $\pm$ 0.089 0.628 $\pm$ 0.080
PUAE 0.576 $\pm$ 0.025 0.745 $\pm$ 0.171 0.800 $\pm$ 0.011 0.613 $\pm$ 0.094
PUSVDD 0.700 $\pm$ 0.021 0.826 $\pm$ 0.137 0.822 $\pm$ 0.016 0.764 $\pm$ 0.044
Table 6: Comparison of anomaly detection performance for unseen anomalies on real datasets.
CIFAR100 PathMNIST OCTMNIST TissueMNIST
IF [Liu et al., 2008] 0.632 $\pm$ 0.005 0.961 $\pm$ 0.055 0.785 $\pm$ 0.004 0.500 $\pm$ 0.187
AE [Hinton and Salakhutdinov, 2006] 0.665 $\pm$ 0.009 0.647 $\pm$ 0.239 0.965 $\pm$ 0.005 0.493 $\pm$ 0.170
DeepSVDD [Ruff et al., 2018] 0.552 $\pm$ 0.033 0.749 $\pm$ 0.200 0.774 $\pm$ 0.072 0.629 $\pm$ 0.069
LOE [Qiu et al., 2022] 0.528 $\pm$ 0.047 0.696 $\pm$ 0.205 0.851 $\pm$ 0.030 0.613 $\pm$ 0.093
ABC [Yamanaka et al., 2019] 0.665 $\pm$ 0.009 0.644 $\pm$ 0.241 0.961 $\pm$ 0.002 0.496 $\pm$ 0.168
DeepSAD [Ruff et al., 2019] 0.561 $\pm$ 0.024 0.755 $\pm$ 0.242 0.874 $\pm$ 0.049 0.651 $\pm$ 0.075
SOEL [Li et al., 2023] 0.570 $\pm$ 0.034 0.776 $\pm$ 0.190 0.918 $\pm$ 0.021 0.666 $\pm$ 0.093
PU [Kiryo et al., 2017] 0.452 $\pm$ 0.039 0.767 $\pm$ 0.215 0.712 $\pm$ 0.154 0.638 $\pm$ 0.098
PUAE 0.671 $\pm$ 0.011 0.806 $\pm$ 0.204 0.941 $\pm$ 0.011 0.574 $\pm$ 0.116
PUSVDD 0.573 $\pm$ 0.027 0.835 $\pm$ 0.215 0.925 $\pm$ 0.022 0.697 $\pm$ 0.114